Two trains approach each other from opposite directions, with one train traveling at 90 mph and and the other at 60 mph. How far apart were the trains originally if took them five hours to pass each other?
Answer by stanbon(75887)

Two trains approach each other from opposite directions, with one train traveling at 90 mph and and the other at 60 mph. How far apart were the trains originally if took them five hours to pass each other?
----
Their apprach rate is 90+60 = 150 mph.
-----
Distance = rate*time = 150*5 = 750 miles
================
Cheers,
Stan H.